OpenStudy (anonymous):

A moving dot P departs from O at an initial speed of 6 m/s in the east direction and accelerates at 2 m/s2 in the east direction. 2 seconds after P's departure, dot Q departs from O to chase down P at a constant speed of k m/s in the east direction. What is the minimum speed, k, required for Q to eventually catch up to P?

Parth (parthkohli):

Brilliant!

OpenStudy (anonymous):

yeah!

hartnn (hartnn):

Let me try, s= ut+1/2 at^2 let Q eventually catch up A at distance 's'. for A, s = 6t1 + 1/2 (2) t1^2 for 'B' the time taken to travel same distance will be t2=t1-2, right ? so, for B, s= k t2 = k(t1-2) so, equating them, k (t1-2) = 6t1 +t1^2 so, k = (6t1 +t1^2)/(t1-2) now to find minimum speed , k you can differentiate that and equate =0

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

let me know if you have further doubts/questions on this.

OpenStudy (anonymous):

wait i'm trying

Parth (parthkohli):

I got 6.

hartnn (hartnn):

^ must be t1, then k=18 is min. speed.

Parth (parthkohli):

OK.

Parth (parthkohli):

thanks

OpenStudy (anonymous):

yes i got 18

hartnn (hartnn):

then its correct.

Parth (parthkohli):

ok